OpenStudy (anonymous):

Someone please help me with matrixes.. I don't know how to do them except on my calculator..

OpenStudy (anonymous):

\[\left[\begin{matrix}-12 & -w ^{2}\\ 2f & 3\end{matrix}\right]=\left[\begin{matrix}2k & -81 \\ -14 & 3\end{matrix}\right]\]

OpenStudy (anonymous):

This is the specific question.. It is asking me to find the variables.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

set them equal is and solve if what you are being asked to do

OpenStudy (misty1212):

so for example \[-12=2k\\ k=-6\]

OpenStudy (misty1212):

\[-w^2=-81\\ w^2=81\\ w=\pm9\]for the second one

OpenStudy (misty1212):

finally solve \[2r=-14\] and you are done

OpenStudy (misty1212):

oops \[2f=-14\] i meant

OpenStudy (anonymous):

okay that makes complete since thank you so much! :) @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]